# 3cm3cm2.3cm2.3cm

### v1

@@ 3.2. Evaluation maps of higher orders
For $\gamma\in{\mathbb{N}}$, there is a Lie superalgebra homomorphism ${\mathfrak{g}}[t]\longrightarrow{\mathfrak{g}}(a,\gamma)$, defined by
@@
[…]
@@
[…]
@@ 3.3. Feigin–Frenkel centers
${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A\otimes t^{-r})=rA\otimes t^{-r-1}$ (3.1)
for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{N}}$. Note that ${\rm T}$ corresponds to the translator operator on $V_{{\rm crit}}({\mathfrak{g}})$, and that ${\mathfrak{z}}(\widehat{{\mathfrak{g}}})$ can be viewed as a commutative subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and is ${\rm T}$-invariant.
@@
$\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\left(A\otimes t^{-1}\right)=-\sum_{i=1}^{\ell}\sum_{k=0}^{\gamma_{i}-1}\frac{A\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}}+\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}$.}$ (3.2)
@@ 3.4. Gaudin algebras for ${\mathfrak{gl}}_{d}$
Let $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ generated by the coefficients of the series $a_{i}(z)$ for $i=1,\ldots,d$. We call $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ the Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$; see (3.2). For $1\leq i\leq\ell$, the singularity at $z_{i}$ is said to be regular if $\gamma_{i}=1$ and irregular otherwise. The commutativity of ${\mathfrak{z}}(\widehat{{\mathfrak{g}}})$ yields the commutativity of $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\left({\mathfrak{z}}(\widehat{{\mathfrak{g}}})\right)$, and hence the algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ is commutative [FFTL, Ry].
@@
The algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ coincides with the subalgebra of $U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ generated by the coefficients of $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\in U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}(\!(z^{-1},\partial_{z}^{-1})\!)$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. This follows from Theorem 3.1 and $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big{(}{\rm T}(S)\big{)}=d/dz\big{(}\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\big{)}$ (cf. [ChL1, Proposition 5.9]).
For any $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$-module $V$, let $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ denote the image of the Gaudin algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ in ${\rm End}(V)$. It is called the Gaudin algebra of $V$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$.
Since $[\partial_{z},z]=1=[-z,\partial_{z}]$, the identity map on the algebra $U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ extends to an isomorphism
${\boldsymbol{\omega}}:U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}(\!(\partial_{z}^{-1},z^{-1})\!)$
@@
^$\widehat{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=-\Big{[}{\boldsymbol{\omega}}\big{(}\delta_{i,j}\partial_{z}-\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big{(}e_{ji}\otimes t^{-1}\big{)}\Big{]}_{i,j=1,\ldots,d}.$ (3.3)
@@
We will use the notations set up in Section 3.3. Recall the symbol $\diamondsuit={p+m|q+n}$ and fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{\diamondsuit}$. Let
@@
for some $b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})$; see [MR, Corollary 3.3] and [ChL1, Section 3.1]. Define $\hat{\mathfrak{z}}_{\diamondsuit}$ to be the subalgebra of $U(t^{-1}{\mathfrak{gl}}_{\diamondsuit}[t^{-1}])$ generated by
@@
Since all $b_{i}$’s belong to ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})$, and ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})$ is ${\rm T}$-invariant, we see that $\hat{\mathfrak{z}}_{\diamondsuit}$ is a commutative subalgebra of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})$.
@@
${\mathcal{L}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big{[}\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big{(}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}\otimes t^{-1}\big{)}\Big{]}_{i,j\in\mathbb{I}}.$ (3.4)
@@
Let $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $U\big{(}{\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ generated by the coefficients of the series $b_{i}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. We call $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ the Gaudin algebra for ${\mathfrak{gl}}_{\diamondsuit}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$. Also, for $1\leq i\leq\ell$, the singularity at $z_{i}$ is said to be regular if $\gamma_{i}=1$ and irregular otherwise.
@@
    We do not know whether Remark 3.2 holds for ${\mathfrak{gl}}_{\diamondsuit}$ in general as it is still a conjecture that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})=\hat{\mathfrak{z}}_{\diamondsuit}$. (Note: The conjecture has been shown to be valid for ${\mathfrak{gl}}_{1|1}$ [MM] and ${\mathfrak{gl}}_{2|1}$ [AN].)
The commutativity of $\hat{\mathfrak{z}}_{\diamondsuit}$ yields the commutativity of $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\left(\hat{\mathfrak{z}}_{\diamondsuit}\right)$. Using an argument similar to the proof of [MR, Corollary 3.6], we see that $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ is a commutative algebra.
For any $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$-module $V$, let $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ denote the image of the Gaudin algebra $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ in ${\rm End}(V)$. We call $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ the Gaudin algebra of $V$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$.
@@ 4.1. Actions on Fock spaces
for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$. The superalgebra $\mathcal{D}$ is naturally a Lie superalgebra with supercommutator [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] defined as in (2.1). Clearly, $\mathcal{D}$ acts naturally on ${\mathcal{F}}$, and so it is a subalgebra of ${\rm End}({\mathcal{F}})$.
@@
$\sum_{i=1}^{p^{\prime}}\gamma_{i}=p,\quad\sum_{i=1}^{q^{\prime}}\gamma_{p^{\prime}+i}=q,\quad\sum_{i=1}^{m^{\prime}}\gamma_{p^{\prime}+q^{\prime}+i}=m,\quad\sum_{i=1}^{n^{\prime}}\gamma_{p^{\prime}+q^{\prime}+m^{\prime}+i}=n.$ (4.1)
@@
${\mathcal{F}}=\left(\bigotimes_{i=1}^{p^{\prime}+q^{\prime}}{\mathbb{C}}\big{[}{\bf y}_{(i)}\big{]}\right)\otimes\left(\bigotimes_{j=p^{\prime}+q^{\prime}+1}^{{\boldsymbol{\ell}}}{\mathbb{C}}\big{[}{\bf x}_{(j)}\big{]}\right)\!.$ (4.2)
@@
${\begin{array}[]{cll}e_{ab}\otimes\overline{t}_{z_{i}}^{k}\mapsto-\sum_{r=l_{i}+1}^{l_{i+1}-k}y^{b}_{r}\partial_{y^{a}_{r+k}},\\ e_{ab}\otimes\overline{t}_{z_{j}}^{l}\mapsto\sum_{r=l_{j}+1-p-q}^{l_{j+1}-l-p-q}(-1)^{|p+q+r|}\partial_{x^{b}_{r}}x^{a}_{r+l},\end{array}}$ (4.3)
for $10\leq k\leq\gamma_{i}-$ 1, $10\leq l\leq\gamma_{j}-$ 1, $1\leq i\leq p^{\prime}+q^{\prime}$, $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$ and $1\leq a,b\leq d$. Thus, $\phi$ extends to a superalgebra homomorphism $\phi:U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}\longrightarrow\mathcal{D}$.
@@
It suffices to verify that $\phi([A,B])=[\phi(A),\phi(B)]$ for all elements $A$ and $B$ on the left-hand side of (4.3). Clearly, for $10\leq k\leq\gamma_{i}-$ 1, $10\leq l\leq\gamma_{j}-$ 1, $1\leq i\leq p^{\prime}+q^{\prime}$, $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$,
@@
Now for $10\leq k\leq\gamma_{i}-$ 1, $10\leq l\leq\gamma_{j}-$ 1, $1\leq i,j\leq p^{\prime}+q^{\prime}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$, we have
@@
We can show similarly that for $10\leq k\leq\gamma_{i}-$ 1, $10\leq l\leq\gamma_{j}-$ 1, $p^{\prime}+q^{\prime}+1\leq i,j\leq{\boldsymbol{\ell}}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$,
@@
${\boldsymbol{\Sigma}}^{(a)}=\left\{\,x^{\alpha}_{i},y^{\alpha}_{r}\;\middle|\;d_{a}+1\leq\alpha\leq d_{a+1},1\leq i\leq m+n,1\leq r\leq p+q\,\right\}.$ (4.4)
@@
${\mathcal{F}}=\bigotimes_{a=1}^{d^{\prime}}{\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}].$ (4.5)
@@
${\begin{array}[]{cll}E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s},\\ E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}},\\ E^{p+q+i}_{s}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}x^{\alpha+k}_{i}y^{\alpha}_{s},\\ E^{p+q+i}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}x^{\alpha+k}_{i}\partial_{x^{\alpha}_{j}},\end{array}}$ (4.6)
for $10\leq k\leq\xi_{a}-$ 1, $1\leq a\leq{d^{\prime}}$, $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$. Thus, $\varphi$ extends to a superalgebra homomorphism $\varphi:U\big{(}{\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big{)}\longrightarrow\mathcal{D}$.
@@
We need to verify that $\varphi([A,B])=\left[\,\varphi(A),\varphi(B)\,\!\right]$ for all elements $A$ and $B$ on the left-hand side of (4.6). Clearly, for $10\leq k\leq\xi_{a}-$ 1, $10\leq l\leq\xi_{b}-$ 1, $1\leq a,b\leq{d^{\prime}}$, $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$,
@@
We will only verify that for $10\leq k\leq\xi_{a}-$ 1, $10\leq l\leq\xi_{b}-$ 1, $1\leq a,b\leq{d^{\prime}}$, $1\leq r,s,s^{\prime}\leq p+q$ and $1\leq i,j\leq m+n$,
$\varphi\!\left(\left[E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k},E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)=\left[\varphi\!\left(E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\right)\!\right]$ (4.7)
@@
$\varphi\!\left(\left[E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k},E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)=\left[\varphi\!\left(E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\right)\!\right]\!.$ (4.8)

### v2

@@ 3.2. Evaluation maps of higher orders
There is a Lie superalgebra homomorphism ${\mathfrak{g}}[t]\longrightarrow{\mathfrak{g}}(a,\gamma)$, defined by
@@
$U({\mathfrak{g}}[t])$$U({\mathfrak{g}}(a,1))$$U({\mathfrak{g}})$$\scriptstyle{{\rm ev}_{a}}$$\scriptstyle{{\rm ev}_{(a,1)}}\scriptstyle{\cong}$
@@
$U({\mathfrak{g}}[t])$$U\left({\mathfrak{g}}\!\left({\bf z},(1^{\ell})\right)\right)$$U({\mathfrak{g}})^{\otimes\ell}$$\scriptstyle{\underline{\rm ev}_{\bf z}}$$\scriptstyle{\underline{{\rm ev}}_{\left({\bf z},(1^{\ell})\right)}}\scriptstyle{\cong}$
@@ 3.3. Feigin–Frenkel centers
(3.1) ${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A\otimes t^{-r})=rA\otimes t^{-r-1}$
[…]
@@
(3.2) $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\left(A\otimes t^{-1}\right)=-\sum_{i=1}^{\ell}\sum_{k=0}^{\gamma_{i}-1}\frac{A\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}}+\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}$.}$
@@ 3.4. Gaudin algebras for ${\mathfrak{gl}}_{d}$
[…]
@@
The algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ coincides with the subalgebra of $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\in U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt]$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. This follows from Theorem 3.1 and $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big({\rm T}(S)\big)=d/dz\big(\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\big)$ (cf. [ChL1, Proposition 5.9]).
For any $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$-module $V$, let $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ denote the image of the Gaudin algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ in ${\rm End}(V)$. It is called the Gaudin algebra of $V$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$.
As in (2.6), the identity map on the algebra $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ extends to an automorphism
${\boldsymbol{\omega}}:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)$
@@
(3.3) $\widehat{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=-\Big[{\boldsymbol{\omega}}\big(\delta_{i,j}\partial_{z}-\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big(e_{ji}\otimes t^{-1}\big)\Big]_{i,j=1,\ldots,d}.$
@@
Recall the symbol $\diamondsuit={p+m|q+n}$ and fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{\diamondsuit}$. Let
@@
[…]
@@
[…]
@@
(3.4) ${\mathcal{L}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big(\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}\otimes t^{-1}\big)\Big]_{i,j\in\mathbb{I}}.$
@@
Let $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of the series $b_{i}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. We call $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ the Gaudin algebra for $\mathfrak{gl}$$_{\diamondsuit}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$. Also, for $1\leq i\leq\ell$, the singularity at $z_{i}$ is said to be regular if $\gamma_{i}=1$ and irregular otherwise.
@@
    We do not know whether Remark 3.2 holds for $\mathfrak{gl}$$_{\diamondsuit}$ in general as it is still a conjecture that $^{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})=\hat{\mathfrak{z}}_{\diamondsuit}$. (Note: The conjecture has been shown to be valid for ${\mathfrak{gl}}_{1|1}$ [$M$] and ${\mathfrak{gl}}_{2|1}$ [AN].)
[…]
For any $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$-module $V$, let $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ denote the image of the Gaudin algebra $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ in ${\rm End}(V)$. We call $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ the Gaudin algebra of $V$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$.
@@ 4.1. Actions on Fock spaces
for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$. The superalgebra $\mathcal{D}$ is naturally a Lie superalgebra with supercommutator $\cdot$$\cdot$] defined as in (2.1). Clearly, $\mathcal{D}$ acts naturally on $\mathcal{F}$}, and so it is a subalgebra of ${\rm End}({\mathcal{F}})$.
@@
(4.1) $\sum_{i=1}^{p^{\prime}}\gamma_{i}=p,\quad\sum_{i=1}^{q^{\prime}}\gamma_{p^{\prime}+i}=q,\quad\sum_{i=1}^{m^{\prime}}\gamma_{p^{\prime}+q^{\prime}+i}=m,\quad\sum_{i=1}^{n^{\prime}}\gamma_{p^{\prime}+q^{\prime}+m^{\prime}+i}=n.$
@@
(4.2) ${\mathcal{F}}=\left(\bigotimes_{i=1}^{p^{\prime}+q^{\prime}}{\mathbb{C}}\big[{\bf y}_{(i)}\big]\right)\otimes\left(\bigotimes_{j=p^{\prime}+q^{\prime}+1}^{{\boldsymbol{\ell}}}{\mathbb{C}}\big[{\bf x}_{(j)}\big]\right)\!.$
@@
(4.3) $\begin{array}[]{cll}e_{ab}\otimes\overline{t}_{z_{i}}^{k}\mapsto-\sum_{r=l_{i}+1}^{l_{i+1}-k}y^{b}_{r}\partial_{y^{a}_{r+k}},\\ e_{ab}\otimes\overline{t}_{z_{j}}^{l}\mapsto\sum_{r=l_{j}+1-p-q}^{l_{j+1}-l-p-q}(-1)^{|p+q+r|}\partial_{x^{b}_{r}}x^{a}_{r+l},\end{array}$}
for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $1\leq i\leq p^{\prime}+q^{\prime}$, $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$ and $1\leq a,b\leq d$. Thus, $\phi$ extends to a superalgebra homomorphism $\phi:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)\longrightarrow\mathcal{D}$.
@@
It suffices to verify that $\phi([A,B])=[\phi(A),\phi(B)]$ for all elements $A$ and $B$ on the left-hand side of (4.3). Clearly, for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $1\leq i\leq p^{\prime}+q^{\prime}$, $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$,
@@
Now for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $1\leq i,j\leq p^{\prime}+q^{\prime}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$, we have
@@
We can show similarly that for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $p^{\prime}+q^{\prime}+1\leq i,j\leq{\boldsymbol{\ell}}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$,
@@
(4.4) ${\boldsymbol{\Sigma}}^{(a)}=\left\{\,x^{\alpha}_{i},y^{\alpha}_{r}\;\middle|\;d_{a}+1\leq\alpha\leq d_{a+1},1\leq i\leq m+n,1\leq r\leq p+q\,\right\}.$
@@
(4.5) ${\mathcal{F}}=\bigotimes_{a=1}^{d^{\prime}}{\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}].$
@@
(4.6) $\begin{array}[]{cll}E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s},\\ E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}},\\ E^{p+q+i}_{s}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}x^{\alpha+k}_{i}y^{\alpha}_{s},\\ E^{p+q+i}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}x^{\alpha+k}_{i}\partial_{x^{\alpha}_{j}},\end{array}$}
for $0\leq k\leq\xi_{a}-1$, $1\leq a\leq{d^{\prime}}$, $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$. Thus, $\varphi$ extends to a superalgebra homomorphism $\varphi:U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)\longrightarrow\mathcal{D}$.
@@
We need to verify that $\varphi([A,B])=\left[\,\varphi(A),\varphi(B)\,\!\right]$ for all elements $A$ and $B$ on the left-hand side of (4.6). Clearly, for $0\leq k\leq\xi_{a}-1$, $0\leq l\leq\xi_{b}-1$, $1\leq a,b\leq{d^{\prime}}$, $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$,
@@
We will only verify that for $0\leq k\leq\xi_{a}-1$, $0\leq l\leq\xi_{b}-1$, $1\leq a,b\leq{d^{\prime}}$, $1\leq r,s,s^{\prime}\leq p+q$ and $1\leq i,j\leq m+n$,
(4.7) $\varphi\!\left(\left[E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k},E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)=\left[\varphi\!\left(E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\right)\!\right]$
@@
(4.8) $\varphi\!\left(\left[E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k},E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)=\left[\varphi\!\left(E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\right)\!\right]\!.$